Carolyn Jean Smith

Home Address 924 Massachusetts Ave. Apt. 2 Arlington, MA 02174 Phone and Email (617) 648-0367 cjsmith@mit.edu

Education  Longy School of Music  1994-present

Master’s candidate in Early Music on recorder. Expected graduation date June 1998.

Workshops Oberlin Baroque Performance Institute, 1993-1997 International Baroque Institute at Longy, 1994-1996

Private lessons Recorder lessons with Sonja Lindblad and Ford Weisberg. Cornetto and voice lessons with Michael Collver. Harpsichord lessons 1996-1997 with Francis Conover Fitch.

Ensembles  Cantata a Trois  founded 1997

Founding member of this chamber music trio, performing primarily cantatas by Telemann and Bach. With Susan Harris, voice, and Ronald Dynneson, harpsichord.

Concerto Scarani founded 1997 Founding member of this ensemble which performs 17th-century Italian canzonas and related music. With Alix Mugar, recorders, Frank Jones, curtal, and Ronald Dynneson, harpsichord and organ.

Stambandet 1995-present Recorder soloist for this Scandinavian vocal ensemble, directed by Allen Levines.

Queens College Collegium Musicum 1982�1983 An 8-12 piece early music ensemble playing mostly Elizabethan music under the direction of Paul Maynard.

Performing Experience  Graduate Recital  April 3, 1997

In partial ful�llment of the requirements for a Master’s Degree in Early Music Performance on Recorder, performed a full-length recital covering music from 1500-1750 in a variety of styles. Recording available upon request.

Recitals MIT Killian Hall 1991, 1992; Longy Wol�nson Room 1989. Recordings available upon request.

Newton Free Library 1995 Performed various works in the Newton Library’s Early Music Sunday concert series.

Longy Early Music Spring 1991 Played recorder in Purcell’s Cupid and Death for the Boston Early Music Festival.

MIT Summer Opera Summer 1990 Played recorder in Handel’s Acis and Galatea under the direction of Steve MacDonald.

Other  Longy Early Dance Workshop  1995-1997

Performed dance music in the context of original dances.

MIT Gilbert & Sullivan Players 1988-1997 Performed in the chorus of many operettas and served in technical positions including producer, house manager, scenic artist, set design, costumes, and board member.